# Leveraging Open-Source Large Language Models for Native Language Identification

### v1

@@ 1 Introduction
Encoder-based transformer models, like BERT Devlin et al. (2019), on the other hand, have yielded poorer performance than conventional ML approaches for the NLI task (Markov et al., 2022; Steinbakken and Gambäck, 2020; Goswami et al., 2024). Previous research suggests that this is likely because NLI concerns very specific linguistic features that models trained on general corpora cannot capture (Markov et al., 2022). Recent research has shown that generative large language models (LLMs) demonstrate promising results for NLI. Lotfi et al. (2020) presented the first study addressing NLI using fine-tuned GPT-2 models, which outperformed previous traditional ML approaches and achieved state-of-the-art results on the NLI benchmark TOEFL11 and ICLE datasets. Zhang and Salle (2023) explored the ability of GPT-3.5 (Brown et al., 2020) and GPT-4 (OpenAI, 2023) to perform NLI. Their results indicate that out-of-the-box GPT models demonstrate outstanding performance, with GPT-4 setting a new performance record of 91.7% accuracy on the TOEFL11 benchmark dataset, and achieve promising results for open-set classification (without a predefined set of L1s), a useful setting for real-world NLI applications.
While Zhang and Salle’s results indicate that LLMs achieve state-of-the-art performance on NLI, they only explore the implementation of GPT-3.5 and GPT-4. The closed-source nature of these models presents a multitude of limitations to research. Providers of closed-source models often disclose minimal information regarding the training data or procedure, hindering the evaluation of results achieved with these models and obscuring biases in training data and models (Balloccu et al., 2024). In addition, closed-source models are typically only accessible via an API, causing lack of control over model updates, which are often communicated poorly to users (Yu et al., 2023; Pozzobon et al., 2023). In turn, the reproducibility of experiments cannot be guaranteed. The usage of closed-source LLMs is also highly costly, which negatively impacts the accessibility of LLMs (Bender et al., 2021).
@@ 2 Data and Models
To comprehensively evaluate the ability of current LLMs to perform NLI, we compare the performance of five open-source LLMs (§2.2), used out-of-the-box and after fine-tuning, with two closed-source commercial LLMs (i.e., GPT-3.5 and GPT-4) on two NLI benchmark datasets.
@@ 2.1 Data
Blanchard et al. (2013): the ETS Corpus of Non-Native Written English (TOEFL11) consists of 12,100 essays, with 1,100 essays per L1, written by English learners with low, medium, or high proficiency. The 11 L1s covered in the data are Arabic (ARA), Chinese (CHI), French (FRE), German (GER), Hindi (HIN), Italian (ITA), Japanese (JPN), Korean (KOR), Spanish (SPA), Telugu (TEL), and Turkish (TUR). We use the TOEFL11 test set for evaluation, which contains 100 essays per L1. The average length of essays in TOEFL11 is 348 words.
@@
2.2 Models
Baselines
Closed-source LLMs
Open-source LLMs
We run inference on the selected open-source LLMs using the same prompt as Zhang and Salle (2023), with the only difference that we instruct each model to respond using JSON dictionaries to restrict the model output to one L1 classification label. For the closed-set task, we include the set of possible L1s in the prompt. If the model classifies an L1 outside of the provided set of classes, we apply iterative prompting up to 5 times. For the open-set task, the prompt does not include a set of possible L1s. For both closed- and open-set tasks, we adapt the prompt to each model’s prompt template. If a prediction cannot be extracted after 5 attempts, the predicted label is set to ‘other’. The prompts for closed-set and open-set tasks are provided in appendices C.1 and C.2, respectively. We use 4-bit quantized instruct versions of the open-source LLMs when prompting out-of-the-box.
3 Results
@@
There are several observations that can be made based on the evaluation results reported in Table 1. Firstly, we observe high accuracy scores on the ICLE-NLI dataset using our implementation of the GPT-3.5 and GPT-4 models. The results are in line with the state-of-the-art results on the TOEFL11 dataset reported in (Zhang and Salle, 2023) and indicate that GPT-4 is able to identify the L1s of highly-proficient English learners both in closed-set and open-set classification experiments.
Secondly, we note a surprisingly low performance of open-source LLMs when used out-of-the-box in a closed-set setting, with the exception of LLaMA-3 on ICLE-NLI. While GPT-4 achieves an accuracy of 91.7% and 95.5% on TOEFL11 and ICLE-NLI, respectively, the five open-source models obtain accuracy scores ranging between 13.6% and 75.8%. All open-source LLMs also perform worse than the baseline approaches, including the simple SVM model with BoW features. Some open-source LLMs tend to predict mostly one or two languages, e.g., Gemma predicting mostly French and LLaMA-2 mostly Chinese, which partially explains such low results.
@@
Previous studies comparing fine-tuned open- and closed-source LLMs provide contradictory findings, with some researchers reporting a drop in accuracy of 16% on sentiment classification for fine-tuned smaller language models (Flan-T5, 770M) compared to ChatGPT (Zhang et al., 2024), while others report that fine-tuned open-source LLMs (Qwen, 7B; LLaMA-3, 8B) outperform closed-source LLMs (GPT-3.5, GPT-4) on text classification tasks (Bucher and Martini, 2024; Edwards and Camacho-Collados, 2024; Wang et al., 2024). The results presented in this study provide evidence that fine-tuned open-source LLMs can achieve comparable performance to closed-source LLMs.
We observe a drop in performance for most open-source LLMs from a closed-set to open-set setting, similarly to closed-source LLMs. Surprisingly, some of the models, i.e., GPT-3.5 and Phi-3, perform better in the open-set than in the closed-set setting. Further research is required to understand the reasons for this behaviour.
@@
Comparing the confusion matrices for GPT-4 and fine-tuned Gemma, the best-performing closed-source and open-source LLM (Appendix B), we note that both models tend to misclassify Hindi texts as Telugu in the TOEFL11 dataset. Hindi and Telugu have been considered a problematic language pair in previous studies on TOEFL11 (Malmasi et al., 2013). Fine-tuned Gemma has a tendency to misclassify Japanese essays as Korean. The high degree of confusion between Korean and Japanese has also been observed in previous research Markov et al. (2022). On ICLE-NLI, GPT-4 erroneously classifies Bulgarian as Russian, both Slavic languages. Gemma misclassifies 14 Czech and Russian samples as Bulgarian. In line with previous research, we note that the confused L1s are either related through geographical location or belong to the same language family.
3.1 Follow-up experiment
In order to verify whether the high performance of GPT-4 is due to potential data contamination, we conducted an additional experiment on an NLI dataset released after the GPT-4 cut-off date: the Varieties of English for Specific Purposes dAtabase (VESPA) [^3] The cut-off date for GPT-4 is September 2021 (OpenAI, 2023), while VESPA was released in 2022. (Paquot et al., 2022). The dataset contains L2 learner English texts written by university students with the following L1s: Dutch (DUT), French (FRE), Spanish (SPA), Norwegian (NOR), and Swedish (SWE). We applied several preprocessing steps to remove XML tags and in-text citations present in the data, and selected a subset of 10 texts per L1 to evaluate GPT-4 performance, 50 texts in total (average essay length is 2,672 words).
We observed a drop of 10–13% on this dataset for GPT-4 compared to its performance on TOEFL11 and ICLE-NLI (82.0% accuracy on a 5-way VESPA subset). Whether this drop is sufficient to indicate that GPT-4 has seen the two benchmark NLI datasets in training requires further analysis, which we leave for future work.
@@ 4 Conclusion
We explored the performance of a variety of open-source LLMs for the NLI task. Our results indicate that open-source LLMs achieve lower performance than closed-source LLMs for this task when used out-of-the-box, while domain-specific fine-tuning of open-source LLMs allows these models to achieve comparable results to the proprietary LLMs, such as GPT-4, on the benchmark TOEFL11 and ICLE-NLI datasets. We believe that our work paves the way for future research on LLM-based Native Language Identification.
@@
More broadly, in our study, we define open-source and closed-source relatively loosely, treating the terms ‘open’ and ‘closed’ as a binary feature to perform a comparative analysis between open-source and closed-source LLMs for NLI. However, there are various dimensions of openness, as a model release involves different components ranging from the disclosure of training datasets to model access (Solaiman, 2023; Liesenfeld and Dingemanse, 2024). Most providers of proclaimed open-source LLMs release little to no information regarding their training data and procedure, despite framing them as being open-source. In turn, it is difficult to determine whether an open-source model’s performance can be attributed to the model’s learning or possible data contamination. The lack of insights into the training data of proclaimed open-source LLMs also hindered our evaluation of LlaMA-3 on ICLE-NLI.
@@ C.2 Open-set prompts
For the open-set experiments, we used the prompt below as an input prompt for all models:
@@ C.3 Fine-tuning prompts
### Instruction: You are a forensic linguistics expert that reads English texts written by non-native authors to classify the native language of the author as one of: “ARA”: Arabic “CHI”: Chinese “FRE”: French “GER”: German “HIN”: Hindi “ITA”: Italian “JPN”: Japanese “KOR”: Korean “SPA”: Spanish “TEL”: Telugu “TUR”: Turkish Use clues such as spelling errors, word choice, syntactic patterns, and grammatical errors to decide on the native language of the author. DO NOT USE ANY OTHER CLASS. IMPORTANT: Do not classify any input as “ENG” (English). English is an invalid choice. Valid output formats: Class: “ARA”, Class: “CHI”, Class: “FRE”, Class: “GER” Classify the text above as one of ARA, CHI, FRE, GER, HIN, ITA, JPN, KOR, SPA, TEL, or TUR. Do not output any other class - do NOT choose “ENG” (English). What is the closest native language of the author of this English text from the given list? ### Input: <<<TOEFL11 ESSAY TEXT>>> ### Response: <<<L1 LABEL>>>

### v2

@@ 1 Introduction
Transformer-based encoder models, like BERT Devlin et al. (2019), on the other hand, have yielded poorer performance than conventional ML approaches for the NLI task (Markov et al., 2022; Steinbakken and Gambäck, 2020; Goswami et al., 2024). Previous research suggests that this is likely because NLI concerns very specific linguistic features that models trained on general corpora cannot capture (Markov et al., 2022). Recent research has shown that generative large language models (LLMs) demonstrate promising results for NLI. Lotfi et al. (2020) presented the first study addressing NLI using fine-tuned GPT-2 models, which outperformed previous traditional ML approaches and achieved state-of-the-art results on the NLI benchmark TOEFL11 and ICLE datasets. Zhang and Salle (2023) explored the ability of GPT-3.5 (Brown et al., 2020) and GPT-4 (OpenAI, 2023) to perform NLI. Their results indicate that out-of-the-box GPT models demonstrate outstanding performance, with GPT-4 setting a new performance record of 91.7% accuracy on the TOEFL11 benchmark dataset, and achieve promising results for open-set classification (without a predefined set of L1s), a useful setting for real-world NLI applications.
While Zhang and Salle’s results indicate that LLMs achieve state-of-the-art performance on NLI, they only evaluate the performance of GPT-3.5 and GPT-4. The closed-source nature of these models presents a multitude of limitations to research. Providers of closed-source models often disclose minimal information regarding the training data or procedure, hindering the evaluation of results achieved with these models and obscuring biases in training data and models (Balloccu et al., 2024). The undisclosed nature of the training data has also raised concerns among researchers about data contamination risks, as it is challenging to determine whether a model’s high performance on a task can be attributed to the model’s effective generalization or potential data leakage (Yu et al., 2023). In addition, closed-source models are typically only accessible via an API, causing lack of control over model updates, which are often communicated poorly to users (Yu et al., 2023; Pozzobon et al., 2023). In turn, the reproducibility of experiments cannot be guaranteed. The usage of closed-source LLMs is also highly costly, which negatively impacts the accessibility of LLMs (Bender et al., 2021).
@@ 2 Data and Models
To comprehensively evaluate the ability of current LLMs to perform NLI, we compare the performance of two closed-source commercial LLMs (i.e., GPT-3.5 and GPT-4) with five open-source LLMs (§2.2), used out-of-the-box and after fine-tuning, on two NLI benchmark datasets.
@@ 2.1 Data
Blanchard et al. (2013): the ETS Corpus of Non-Native Written English (TOEFL11) consists of 12,100 essays, with 1,100 essays per L1, written by English learners with low, medium, or high proficiency levels. The 11 L1s covered in the data are Arabic (ARA), Chinese (CHI), French (FRE), German (GER), Hindi (HIN), Italian (ITA), Japanese (JPN), Korean (KOR), Spanish (SPA), Telugu (TEL), and Turkish (TUR). We use the TOEFL11 test set for evaluation, which contains 100 essays per L1. The average length of essays in TOEFL11 is 348 words.
@@
2.2 Models
Baselines
Closed-source LLMs
Open-source LLMs
We run inference on the selected open-source LLMs using the same prompt as Zhang and Salle (2023), with the only difference that we instruct each model to respond using JSON dictionaries to restrict the model output to one L1 classification label. For the closed-set task, we include the set of possible L1s in the prompt. If the model classifies an L1 outside of the provided set of classes, we apply iterative prompting up to 5 times. For the open-set task, the prompt does not include a set of possible L1s. For both closed- and open-set tasks, we adapt the prompt to each model’s prompt template. If a prediction cannot be extracted after 5 attempts, the predicted label is set to ‘other’. The prompts for closed-set and open-set tasks are provided in appendices C.1 and C.2, respectively. We use 4-bit quantized instruction-fine-tuned versions of the open-source LLMs when prompting out-of-the-box.
3 Results
3.1 Closed-Source LLMs
We observe high accuracy scores on the ICLE-NLI dataset in our experiments using the GPT-3.5 and GPT-4 models. The results are in line with the state-of-the-art results on the TOEFL11 dataset reported in (Zhang and Salle, 2023) and indicate that GPT-4 is able to identify the L1s of highly-proficient English learners both in closed-set and open-set classification experiments.
3.2 Open-Source LLMs Out-of-the-Box
We note a surprisingly low performance of open-source LLMs when used out-of-the-box in a closed-set setting, with the exception of LLaMA-3 on ICLE-NLI. While GPT-4 achieves an accuracy of 91.7% and 95.5% on TOEFL11 and ICLE-NLI, respectively, the five open-source models obtain accuracy scores ranging between 13.6% and 75.8%. All open-source LLMs also perform worse than the baseline approaches, including the simple SVM model with BoW features. Some open-source LLMs tend to predict mostly one or two languages, e.g., Gemma predicting mostly French and LLaMA-2 mostly Chinese, which partially explains such low results. The large performance gap raises the concern that closed-source LLMs might have seen the NLI benchmark datasets in training. Additional research is required to explore the possibility of data leakage, e.g., by examining whether a model has memorized a given text using perplexity measurements (carlini_extracting-2021).
3.3 Fine-Tuned Open-Source LLMs vs. Closed-Source LLMs
@@ 3.2 Open-Source LLMs Out-of-the-Box
Previous studies comparing closed-source and fine-tuned open-source LLMs provide contradictory findings, with some researchers reporting a drop in accuracy of 16% on sentiment classification for fine-tuned smaller language models (Flan-T5, 770M) compared to ChatGPT (Zhang et al., 2024), while others report that fine-tuned open-source LLMs (Qwen, 7B; LLaMA-3, 8B) outperform closed-source LLMs (GPT-3.5, GPT-4) on text classification tasks (Bucher and Martini, 2024; Edwards and Camacho-Collados, 2024; Wang et al., 2024). The results presented in this study provide evidence that fine-tuned open-source LLMs can achieve comparable performance to closed-source LLMs.
@@
Comparing the confusion matrices for GPT-4 and fine-tuned Gemma, the best-performing closed-source and open-source LLMs (Appendix B), we note that both models tend to misclassify Hindi texts as Telugu in the TOEFL11 dataset. Hindi and Telugu have been considered a problematic language pair in previous studies on TOEFL11 (Malmasi et al., 2013). Fine-tuned Gemma has a tendency to misclassify Japanese essays as Korean. The high degree of confusion between Korean and Japanese has also been observed in previous research Markov et al. (2022). On ICLE-NLI, GPT-4 erroneously classifies Bulgarian as Russian, both Slavic languages. Gemma misclassifies 14 Czech and Russian samples as Bulgarian. In line with previous research, we note that the confused L1s are either related through geographical location or belong to the same language family.
3.4 Closed-Set and Open-Set Settings
We observe a drop in performance for most open-source LLMs from a closed-set to open-set setting, similarly to closed-source LLMs. Surprisingly, some of the models, i.e., GPT-3.5 and Phi-3, perform better in the open-set than in the closed-set setup. Further research is required to understand the reasons for this behaviour.
@@ 4 Conclusion
We explored the performance of a variety of open-source LLMs for the NLI task. Our results indicate that open-source LLMs achieve lower performance than closed-source LLMs for this task when used out-of-the-box, while domain-specific fine-tuning of open-source LLMs allows these models to achieve comparable results to the proprietary LLMs, such as GPT-4, on the benchmark TOEFL11 and ICLE-NLI datasets. We believe that our work opens up avenues for future research on LLM-based Native Language Identification. Future research could explore few-shot prompting and different prompt variations as a way to potentially boost the performance of open-source LLMs.
@@
More broadly, in our study, we define open-source and closed-source relatively loosely, treating the terms ‘open’ and ‘closed’ as a binary feature to perform a comparative analysis between open-source and closed-source LLMs for NLI. However, there are various dimensions of openness, as a model release involves different components ranging from the disclosure of training datasets to model access (Solaiman, 2023; Liesenfeld and Dingemanse, 2024). Most providers of proclaimed open-source LLMs release little to no information regarding their training data and procedure, despite framing them as being open-source. In turn, it is difficult to determine whether an open-source model’s performance can be attributed to the model’s learning or possible data contamination. The lack of insights into the training data of proclaimed open-source LLMs also hindered our evaluation of LlaMA-3 on the ICLE-NLI dataset.
@@ C.2 Open-Set Prompts
For the open-set experiments, we used the prompt below as an input prompt for all the models:
@@ C.3 Fine-Tuning Prompts
### Instruction: You are a forensic linguistics expert that reads English texts written by non-native authors to classify the native language of the author as one of: “ARA”: Arabic “CHI”: Chinese “FRE”: French “GER”: German “HIN”: Hindi “ITA”: Italian “JPN”: Japanese “KOR”: Korean “SPA”: Spanish “TEL”: Telugu “TUR”: Turkish Use clues such as spelling errors, word choice, syntactic patterns, and grammatical errors to decide on the native language of the author. DO NOT USE ANY OTHER CLASS. IMPORTANT: Do not classify any input as “ENG” (English). English is an invalid choice. Valid output formats: Class: “ARA”, Class: “CHI”, Class: “FRE”, Class: “GER” Classify the text below as one of ARA, CHI, FRE, GER, HIN, ITA, JPN, KOR, SPA, TEL, or TUR. Do not output any other class - do NOT choose “ENG” (English). What is the closest native language of the author of this English text from the given list? ### Input: <<<TOEFL11 ESSAY TEXT>>> ### Response: <<<L1 LABEL>>>